An ion source is producing 6Li ions, which have a charge +e and mass 9.99 x 10-27 kg. The ions are accelerated by a potential difference of 10.0 kV and pas horizontally into a region in which there is uniform vertical magnetic field of magnitude B = 1.78 T. Calculate the strength of the smallest electric field to be set up over the same region, that will allow the 6Li ions to pass through undeflected. Express your answer in units of 10-6 N/C (Specify your answer up to two decimal places)
Answer #1

By Conservation of energy

(1/2)mv2 = qV

(1/2)(9.99*10-27)v2=(1.6*10-19)(10*103)

v=5.66*106 m/s

electric field is given by

E=vB=(5.66*106 )(1.78)

E=1.01*106 V/m